OpenStudy (anonymous):

can someone please show me how tofind the integral of x^3sinx^2dx

OpenStudy (anonymous):

u use the integration by parts formula INT (u dv) = uv - INT vdu

OpenStudy (anonymous):

yah i know that but im actually having trouble setting it up

OpenStudy (anonymous):

ok - u want to make INTvdu easy to do if possible try u = x^3 and dv = sin^2x and if this is problem reverse them

OpenStudy (anonymous):

for some reason people first u substitution bu i do no know why

OpenStudy (anonymous):

there is a recomendation in the literature but i cant remember it - I just dive in and see what happens!